OpenStudy (anonymous):

Determine whether the given equation is linear or nonlinear: 2x-7=πx-2

OpenStudy (unklerhaukus):

\[2x-7=πx-2\]\[(2-π)x=5\]\[x={5 \over 2-π}\] the equation is a vertical line